Night report 2021-08-04

Observers: Priscila Pessi, Lluís Galbany (night plan)

Support:

TAT: Nada Ihanec, Emmy Paraskeva, Claudia Gutiérrez, Jacob Teffs, Michael Fulton, Janet Chen

EFOSC Calibrations:

Bias, domeflats

Weather:

Start: humidity ~ 29%, wind ~ 7.3m/s, seeing ~ 1.1"

Middle: humidity ~ %, wind ~ m/s, seeing ~ "

End: humidity ~ 25%, wind ~ 3.9m/s, seeing ~ 1.5"

Comments (times in UT):

There was a power outage at La Silla which means all systems shut down.

The recovery process took a couple of ours. After opening some time was needed to stabilize telescope temperature.

Once the telescope was stabilized image analysis was performed and observations started.